# Train Once, Deploy Anywhere: Realize Data-Efficient Dynamic Object Manipulation

### v1

@@ V-B Method Comparison
This subsection compares GEM with previous manipulation methods in dynamic object manipulation. Although the classic kinematics-based methods can also handle dynamic object manipulation after precise system identification and careful parameter tuning, they demand special algorithm designs for different tasks and testing environments. This task-dependent characteristic violates the focus of this work, how to improve the generalization of a manipulation system. Therefore, we mainly compare GEM with the recent learning-based methods, which present better generalization.
The compared methods include AnyGrasp [37], ConvMLP[62], RVT-2 [63], ACT [8], Diffusion Policy (DP) [46], OpenVLA [11], and PI0 [12]. Among them, AnyGrasp can generate grasp pose candidates based on the 3D points of the target object. Then, the robot arm selects the grasp pose with the highest confidence and grasps the target based on kinematic control. Since AnyGrasp only predicts the grasp pose of each object for once, the grasp is executed in an open-loop manner. RVT-2 is a representative two-stage method that takes RGB-D information as input and processes this information based on point cloud reconstruction. The inputs of the other compared methods are the wrist and external view images. Among them, ConvMLP, ACT, and DP are lightweight neural networks based on image observations. OpenVLA and PI0 are two influential vision-language-action (VLA) models. As OpenVLA and PI0 contain billions of parameters, their inference speeds are limited.
@@
AnyGrasp 0.41 - - - - 0.43 - - - - - 4.35
ConvMLP 0.06 0.19 0.05 0.00 0.08 0.00 0.08 0.00 0.00 0.02 0.05 17.54
ConvMLP* 0.28 0.48 0.19 0.01 0.24 0.06 0.35 0.02 0.00 0.11 0.18 13.16
RVT-2 0.08 0.24 0.03 0.00 0.09 0.00 0.14 0.00 0.00 0.04 0.07 31.48
RVT-2* 0.47 0.62 0.15 0.27 0.38 0.18 0.50 0.00 0.12 0.20 0.29 19.69
ACT 0.14 0.29 0.10 0.00 0.13 0.05 0.22 0.00 0.00 0.07 0.10 43.48
ACT* 0.62 0.75 0.53 0.44 0.59 0.26 0.65 0.08 0.27 0.32 0.46 23.70
DP 0.09 0.24 0.07 0.00 0.10 0.00 0.17 0.00 0.00 0.04 0.07 27.32
DP* 0.55 0.70 0.46 0.36 0.52 0.18 0.54 0.05 0.23 0.25 0.39 17.86
OpenVLA 0.11 0.26 0.08 0.00 0.11 0.07 0.24 0.01 0.00 0.08 0.10 1.61
OpenVLA* 0.64 0.74 0.46 0.35 0.55 0.35 0.68 0.06 0.29 0.35 0.45 1.58
PI0 0.23 0.36 0.18 0.00 0.19 0.12 0.29 0.10 0.00 0.13 0.16 1.97
PI0* 0.68 0.81 0.59 0.65 0.68 0.50 0.75 0.14 0.43 0.46 0.57 1.92
GEM (Ours) 0.88 0.98 0.83 0.80 0.87 0.77 0.91 0.70 0.72 0.78 0.83 23.81
TABLE II: Success rate with previous methods in real world tasks.
@@
AnyGrasp 0.12 - - - -
ConvMLP 0.00 0.00 0.00 0.00 0.00
ConvMLP* 0.06 0.38 0.02 0.00 0.12
RVT-2 0.02 0.15 0.00 0.00 0.04
RVT-2* 0.17 0.59 0.08 0.04 0.22
ACT 0.08 0.23 0.05 0.00 0.09
ACT* 0.53 0.76 0.40 0.35 0.51
DP 0.05 0.17 0.01 0.00 0.06
DP* 0.46 0.70 0.32 0.25 0.43
OpenVLA 0.00 0.00 0.00 0.00 0.00
OpenVLA* 0.29 0.35 0.14 0.00 0.20
PI0 0.00 0.00 0.00 0.00 0.00
PI0* 0.37 0.58 0.21 0.00 0.29
GEM (Ours) 0.92 0.96 0.79 0.74 0.85
@@ V-B1 Simulated Comparison
This part compares GEM with the aforementioned counterparts in the simulation tasks described in Section V-A. Since AnyGrasp is for generating grasp poses, we only test it in Pick tasks. The algorithms in the hybrid action control module are utilized to help AnyGrasp handle the moving dynamics of objects. The other compared methods are trained using the same demonstrations as our recurrent probabilistic network. Because the compared models have no designs for dynamic object manipulation, we test them in two modes, the sole mode and hybrid mode. The sole mode means the robot movement is solely controlled by a network output. In the hybrid mode, the network output is combined with visual servo tracking like our developed hybrid action control module. In this way, the original dynamic object manipulation problem is converted into a static problem, which favors the performances of the compared methods.
As mentioned in Section IV-F, we collect 30 diverse assets to generate demonstrations and train the compared methods except AnyGrasp. To verify their out-of-domain generalization on unseen geometries, we additionally select 14 new assets, as shown in Fig. 10, to conduct experiments. The out-of-domain Insert validation is conducted using blocks similar to the ones shown in Fig. 9 but we randomly change the block sizes.
@@
The success rates and inference speeds of all the methods are presented in Table I. Specifically, the $2_{\rm nd}-6_{\rm th}$ columns are the success rates of in-domain evaluations based on the assets shown in Fig. 9 (b). The $7_{\rm th}-11_{\rm th}$ columns correspond to the out-of-domain success rates of using the assets in Fig. 10. Avg_in, Avg_out, and Avg refer to the average success rates of in-domain tests, out-of-domain tests, and all tests, respectively. The inference speeds reported in the $13_{\rm th}$ column are the model inference frequencies tested on an RTX4090 GPU. An asterisk (*) following the method name indicates that the method is validated in the hybrid mode. The hybrid mode increases the inference time compared to the sole mode (denoted without the asterisk) because of the additional adoption of an instance segmentation model.
@@
According to the results, we can find that GEM achieves the best success rates in performing all the skills. This is partly because of the efficient structure of the recurrent probabilistic network on perceiving object geometries based on points. In addition, it is observed that the methods in the hybrid mode significantly outperform the ones in the sole mode, which is attributed to that our hybrid action control module alleviates the manipulation challenge. For AnyGrasp, we find it generates promising pick poses. However, it works in an open-loop manner, which restricts its effectiveness. Though RVT-2 uses depth information like our method, its network architecture is inefficient in geometry information exploitation, resulting in its unsatisfactory success rates. Comparing the results of the VLA models and other lightweight networks like ACT, we observe VLAs generally present better out-of-domain generalization thanks to they are pre-trained on numerous data. However, the speeds of VLAs are slower. The slow inference does not affect the success rates significantly in simulated environments because the simulation is paused during model inference, but it harms performances in real tasks significantly.
@@ V-B2 Real Robot Comparison
This study compares GEM with the aforementioned counterparts in real-world tasks using the experiment platform introduced in Fig 5. The objects illustrated in Fig. 11, which are in diverse geometries and sizes, are selected to conduct this comparison. Specifically, we use the left nine objects in Fig. 11 to collect 20 demonstrations for every skill. Then, the total of 80 demonstrations are utilized to fine-tune the policies that have been trained using the simulation data. Finally, employing the same nine objects, we test their in-domain performances in the hybrid mode.
We do not compare the out-of-domain results in this experiment because we find the generalization of the compared methods is generally limited. This is because the color differences between in-domain and out-of-domain objects disturb their action prediction. In the simulation, all the objects are rendered with pure colors, so the compared methods generalize successfully to some extent according to the results in Table I. Nevertheless, in real-world validation, the appearances between in-domain and out-of-domain objects are quite different. Differently, GEM can directly generalize to real world-tasks even though trained utilizing only simulated data, according to the sim-to-real results in Section V-E. This result highlights the advantage of using uncolored 3D points.
@@
The real robot comparison result is presented in Table II, where we can find that GEM presents high success rates after tuning using only 80 demonstrations. Besides, the results of the compared methods in the hybrid mode significantly outperform the results in the sole mode, indicating the value of combining classic control and learning algorithms. Moreover, we can find that although PI0 behaves better than ACT in simulated environments, its success rates are lower in real robot experiments. This is caused by the slow inference of PI0. Therefore, fast inference speed is important for models used in dynamic object manipulation.
Figure 12: Subfigure (a) illustrates the MazeNav task, where a robot is initialized in a random grid of the $5\times 5$ maze and needs to find a path to reach the yellow star. The yellow star is always in the botton-right corner of the maze. Subfigure (b) visualizes the action of the expert policy in every grid.
V-C Theory Verification
This subsection validates the entropy analysis framework described in Section III, i.e., how the changes of observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and action prediction entropy $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ affect the target distribution approximation precision. As the entropy of robotic manipulation is too complex to quantify, we design a toy task named MazeNav to study this problem. Specifically, as shown in Fig. 12 (a), a robot is initialized in a random grid of the $5\times 5$ maze and needs to find a path to reach the destination marked by a yellow star. The destination is always in the bottom-right corner of the maze. The action space includes moving up, down, left, and right. The robot is not allowed to move beyond the maze boundaries.
As illustrated in Fig. 12 (b), a deterministic expert policy is built to generate demonstrations. Based on the demonstrations, a student policy is to estimate the expert policy action distribution by calculating the frequency of each action taken in each state observation. If an action is never performed given a state, the student policy assigns equal possibilities to all candidate actions in this observation state.
Besides the robot position, the state observation of the student policy includes an integer variable $n_{m}$ randomly sampled from $1,N_{m}]$. As a random variable, $n_{m}$ is irrelevant to action prediction. We can change $N_{m}$ to adjust $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. To alter $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, we introduce a threshold variable $\eta\in(0,1)$. During collecting demonstrations, if a variable sampled between $(0,1)$ is smaller than $\eta$, the action is randomly selected from the action space rather than decided by the expert policy. Hence, by increasing $\eta$, $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is enhanced.
Figure 13: This figure illustrates the relation between $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and the target distribution approximation precision. $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is adjusted by changing $N_{m}$. The approximation precision is measured by the KL divergence between the expert distribution and student distribution.
Figure 14: This figure illustrates the relation between $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ and the target distribution approximation precision. $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is adjusted by changing $\eta$. The approximation precision is measured by the KL divergence between the expert distribution and student distribution.
V-C1 Observation Entropy Verification
By adjusting $N_{m}$ and the demonstration trajectory number, this part studies the relation between $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and the expert distribution approximation precision. Specifically, $\eta$ is fixed as 0, meaning that there is no action prediction ambiguity. To study the approximation performance under various $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ values, we repeat this experiment by setting $N_{m}$ to 1, 3, and 5 separately. In each experiment, the expert policy generates 10, 20, 30, 40, and 50 demonstration trajectories to update the student policy, respectively. After the update, the expert distribution approximation precision is measured by the KL divergence between the expert and student action distributions.
To eliminate the influence of random factors, we set different seeds and repeat each experiment five times to calculate the standard deviation. The obtained KL divergence values and standard deviations under different $N_{m}$ and demonstration trajectory numbers are visualized in Fig. 13. We can observe that increasing the demonstration number improves the approximation performance and is helpful in reducing the standard deviation. In addition, by increasing $N_{m}$, which means larger $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, the approximation precision deteriorates when using the same number of demonstrations. These observations are consistent with the theoretical analysis in Section III-C.
V-C2 Action Prediction Entropy Verification
This part explores the relation between $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ and the expert distribution approximation precision. Specifically, $N_{m}$ is fixed as 1, indicating that the policy observation does not include redundant information. We set $\eta$ to 0.0, 0.3, and 0.6 separately to adjust $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$. Similar to the previous experiment, we update the policy using 10, 20, 30, 40, and 50 demonstration trajectories, respectively. The experiment of each hyper-parameter is repeated five times to compute the mean KL divergence and standard deviation. The experimental results of different $\eta$ and demonstration numbers are shown in Fig. 14.
According to the results, we can find that a larger $\eta$, which indicates bigger $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, leads to poorer distribution estimation precision. In addition, by incorporating more demonstrations, the student policy can still estimate a promising expert policy distribution even though $\eta$ is high. These results confirm the theoretical analysis in Section III-D.
Figure 15: This experiment suggests that the student policy can perfectly estimate the action choices of the expert policy if sufficient demonstrations are provided, even though the demonstrations are very noisy (high $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$) and contain much irrelevant observation (high $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$). Specifically, the subfigure (a) visualizes the expert policy. Subfigures (b)-(d) show the student policies updated with 50, 200, and 1000 demonstrations, respectively. The student policy using 1000 demonstrations perfectly replicates the expert policy.
V-C3 Effect of Data Scaling Up
This experiment aims to validate the hypothesis in Section III-B that the optimal actions can be precisely approximated by scaling up expert demonstrations. To study this hypothesis, we set $N_{m}$ to 10, which leads to high $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. Besides, $\eta$ is 0.9, meaning the robot performs random actions in 90% of the steps. This results in a large $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$.
We update the student policy with 50, 200, and 1000 demonstrations. Then, we visualize the student policy actions corresponding to the biggest probabilities under various states in Fig. 15 (b)-(d). We can find that even though the demonstration data is highly noisy, the student policy still perfectly replicates the expert policy using 1000 demonstrations. This result confirms the hypothesis that the optimal actions can be precisely approximated with sufficient demonstrations.
TABLE III: Comparison between point and imgae observations.
@@ V-C3 Effect of Data Scaling Up
Image $\times$ 0.75 0.60 0.64 0.33
Image $\checkmark$ 0.41 0.28 0.38 0.14
Point $\times$ 0.79 0.71 0.73 0.66
Point $\checkmark$ 0.87 0.78 0.85 0.74
Figure 16: This figure visualizes some instance segmentation cases during manipulation, where we can find that the segmentation masks are precise.
V-D Component Analysis
@@
This subsection ablates the various designs in GEM. Due to the space limit, when simultaneously reporting the in-domain and out-of-domain results in simulated and real environments, we report the average success rates of all skills.
V-D1 Analysis on Point based Representation
Different from previous works commonly taking 2D images as input, our method utilizes 3D points. This design is based on the intuition that accurate manipulation relies on the geometry of objects, and 3D points better deliver such information. This experiment is to validate this intuition. The results are presented in Table III, where the first two rows of results correspond to the models based on images, and the last two rows are the models using points. When the input is in the image modality, the feature extractor of the recurrent probabilistic network is changed to DINOv2 [64], a widely adopted network due to its discriminative representation. The second column indicates whether the background regions are removed based on instance segmentation results. “Sim_in”, “Sim_out”, “Real_in”, and “Real_out” refer to the average success rates in the settings of in-domain simulation, out-of-domain simulation, in-domain real world, out-of-domain real world, respectively.
According to the results, we can first find that the models taking 3D points as input outperform the ones with images, which confirms our intuition. Secondly, it is observed that the operation of removing background regions favors point based models but harms the performance of image based models. This is because when the input is an image, the model demands background texture to infer the depth of the target object. By contrast, the model does not need background since the object position has been known from 3D points.
V-D2 Analysis on Instance Segmentation
The results in Table III have revealed how the operation of removing background based on instance segmentation affects manipulation success rates. This part analyzes the efficiency and robustness of our employed instance segmentation model, YOLOv11. Since the official YOLOv11 only supports segmenting objects in the categories defined in the COCO dataset [65], we need to collect a few additional instance segmentation data to enable YOLOv11 to handle our concerned categories. Thankfully, the development of SAM2 [57] makes this data collection process efficient. Specifically, based on SAM2, we only need to record a video for every concerned category of objects and click a point on each object to generate segmentation masks. In this way, obtaining tuning data totally takes several minutes. The inference speed of YOLOv11 evaluated on an RTX4090 GPU is more than 100FPS, meeting the real-time requirement.
Afterwards, to evaluate the segmentation precision of the tuned YOLOv11, we collect a dataset consisting of 1800 images for the objects shown in Fig. 11 (200 frames per object). Then, we utilize this dataset to validate the segmentation precision of YOLOv11. Following the testing protocol in COCO, YOLOv11 achieves a Mask AP of 94.5% in this dataset, suggesting that the segmentation precision YOLOv11 satisfies the manipulation task demand. Furthermore, we visualize some challenging cases in Fig. 16 during manipulation. It is found that the segmentation quality is promising.
TABLE IV: Manipulation performance in various moving speeds.

### v2

@@ V-B Method Comparison
This subsection compares GEM with previous manipulation methods in dynamic object manipulation. Although the classic kinematics-based methods can also handle dynamic object manipulation after precise system identification and careful parameter tuning, they demand special algorithm designs for different tasks and testing environments. This task-dependent characteristic violates the focus of this work, how to build a policy that is generalizable across various scenarios. Therefore, we mainly compare GEM with the recent learning-based methods, which present better generalization.
The compared methods include AnyGrasp [12], ConvMLP[57], RVT-2 [14], ACT [60], Diffusion Policy (DP) [7], Diffusion Policy 3D (DP3) [56], OpenVLA [24], and PI0 [4]. Among them, AnyGrasp can generate grasp pose candidates based on the 3D points of the target object. Then, the robot arm selects the grasp pose with the highest confidence and grasps the target based on kinematic control. Since AnyGrasp only predicts the grasp pose of each object once, the grasp is executed in an open-loop manner. RVT-2 is a representative two-stage method that takes RGB-D information as input and processes this information based on point cloud reconstruction. Similar to the GEM network, DP3 takes 3D points as environmental observation, but it does not employ visual color feature. The inputs of the other compared methods are the wrist and external view images. Among them, ConvMLP, ACT, and DP are lightweight neural networks based on image observations. OpenVLA and PI0 are two influential vision-language-action (VLA) models. As OpenVLA and PI0 contain billions of parameters, their inference speeds are limited.
@@
AnyGrasp* 0.41 - - - - 0.46 - - - - - 4.35
ConvMLP 0.06 0.19 0.05 0.00 0.08 0.00 0.03 0.00 0.00 0.01 0.05 17.54
ConvMLP* 0.28 0.48 0.19 0.01 0.24 0.03 0.24 0.02 0.00 0.08 0.16 13.16
RVT-2 0.08 0.24 0.03 0.00 0.09 0.00 0.07 0.00 0.00 0.02 0.05 31.48
RVT-2* 0.47 0.62 0.15 0.27 0.38 0.12 0.40 0.00 0.05 0.14 0.26 19.69
ACT 0.14 0.29 0.10 0.00 0.13 0.02 0.19 0.00 0.00 0.05 0.09 43.48
ACT* 0.62 0.75 0.53 0.44 0.59 0.20 0.39 0.07 0.23 0.22 0.41 23.70
DP 0.09 0.24 0.07 0.00 0.10 0.00 0.15 0.00 0.00 0.04 0.07 27.32
DP* 0.55 0.70 0.46 0.36 0.52 0.17 0.32 0.03 0.18 0.18 0.35 17.86
DP3 0.31 0.50 0.29 0.00 0.28 0.19 0.40 0.22 0.00 0.20 0.24 29.40
DP3* 0.72 0.85 0.77 0.70 0.76 0.57 0.74 0.62 0.44 0.59 0.68 18.87
OpenVLA 0.11 0.26 0.08 0.00 0.11 0.06 0.18 0.01 0.00 0.06 0.09 1.61
OpenVLA* 0.64 0.74 0.46 0.35 0.55 0.31 0.64 0.05 0.24 0.31 0.43 1.58
PI0 0.23 0.36 0.18 0.00 0.19 0.09 0.24 0.06 0.00 0.10 0.15 1.97
PI0* 0.68 0.81 0.59 0.65 0.68 0.47 0.70 0.11 0.38 0.42 0.55 1.92
GEM (Ours) 0.94 1.00 0.90 0.86 0.93 0.86 0.96 0.90 0.83 0.89 0.91 23.81
TABLE II: Success rate comparison with previous methods in real world tasks.
@@
AnyGrasp* 0.45 - - - -
ConvMLP 0.08 0.25 0.07 0.00 0.10
ConvMLP* 0.36 0.55 0.28 0.13 0.33
RVT-2 0.13 0.30 0.09 0.00 0.13
RVT-2* 0.51 0.68 0.41 0.28 0.47
ACT 0.18 0.34 0.20 0.00 0.18
ACT* 0.65 0.76 0.55 0.49 0.61
DP 0.15 0.29 0.13 0.00 0.14
DP* 0.61 0.69 0.56 0.32 0.55
DP3 0.36 0.45 0.39 0.04 0.31
DP3* 0.70 0.80 0.72 0.61 0.71
OpenVLA 0.00 0.02 0.05 0.00 0.02
OpenVLA* 0.32 0.39 0.20 0.03 0.24
PI0 0.03 0.08 0.14 0.00 0.06
PI0* 0.40 0.48 0.56 0.11 0.39
GEM (Ours) 0.98 1.00 0.93 0.89 0.95
@@ V-B1 Simulated Comparison
This part compares GEM with the aforementioned counterparts in simulation. Since AnyGrasp is for generating grasp poses, we only test it in Pick tasks. The other compared methods are trained using the same simulated demonstrations as our GEM network. Since the compared models have no designs for dynamic object manipulation, we test them in two modes, the sole mode and hybrid mode. The sole mode means the robot movement is solely controlled by their network outputs. In the hybrid mode, we utilize a segmentation model [21] to estimate the spatial position of the target object and then employ a visual servo control algorithm [38] that controls the end-effector to track the target object. The control signal is the sum of the network output and the visual servo control module, which shares a similar process as our proposed manipulation action decomposition strategy. In this way, the original dynamic object manipulation problem is converted into a static object manipulation problem, significantly improving the performance of the compared methods. As AnyGrasp is an open-loop method, we only report its results in the hybrid mode.
@@
As described in Section V-A, we compare these policies in the in-domain setting (manipulate seen objects in Isaac Gym) and out-of-domain setting (manipulate unseen objects in Isaac Sim). The success rates and inference speeds of all the methods are presented in Table I. Specifically, the $2_{\rm nd}-6_{\rm th}$ columns are the success rates of the in-domain setting. The $7_{\rm th}-11_{\rm th}$ columns correspond to the out-of-domain setting. Avg_in, Avg_out, and Avg refer to the average success rates of in-domain tests, out-of-domain tests, and all tests, respectively. The inference speeds reported in the $13_{\rm th}$ column are the model inference frequencies tested on an RTX4090 GPU. An asterisk (*) following a method name indicates that the method is validated in the hybrid mode. The hybrid mode increases the inference time compared to the sole mode (denoted without the asterisk) because of the additional adoption of an instance segmentation model.
@@
According to the results, we can find that GEM achieves the best results in performing all the skills. This is partly because of the design that helps the GEM network prioritize geometry information. In addition, it is observed that the methods tested in the hybrid mode significantly outperform those of the sole mode, which suggests the effectiveness of our action decomposition strategy. For AnyGrasp, we find it generates promising pick poses. However, it works in an open-loop manner, which restricts its real-time response to object pose changes. Though RVT-2 uses depth information like our method, its network architecture is inefficient in geometry information exploitation, resulting in its unsatisfactory success rates. Comparing the results of the VLA models and other lightweight networks like ACT, we observe that VLAs generally present better out-of-domain generalization because they are pre-trained on numerous data. However, the speeds of VLAs are slower. The slow inference does not affect the success rates significantly in simulated environments because the simulation is paused during model inference, but it harms performance in real-world tasks significantly. Compared with the other policies, the success rates of DP3 are promising although its structure is simple. This is because it takes 3D points without visual texture as input, and this strategy also prioritizes 3D geometry.
@@ V-B2 Real Robot Comparison
This study compares GEM with the aforementioned counterparts in real-robot tasks using the platform shown in Fig. 8 (a). We only report the results in the tuned sim-to-real setting, which employs 10 real-world demonstrations to fine-tune policies before evaluation. This is because we find the compared methods mostly cannot generalize to real-world manipulation after training with simulated data, though GEM generalizes well since it prioritizes geometry information. The tuning step adapts the policy to the real-world embodiment and environment observations.
@@
Similar to the simulation experiments, we validate the compared methods in both the sole and hybrid modes. The real robot comparison results are presented in Table II, where we can find that GEM presents high success rates. According to the results, the sim-to-real performance of GEM outperforms the compared counterparts by large margins. Besides, the results of the compared methods in the hybrid mode significantly outperform the results in the sole mode, indicating the value of our manipulation action decomposition. Moreover, we can find that although PI0 behaves better than ACT in simulated environments, its success rates are lower in real robot experiments. This is caused by the slow inference of PI0. Therefore, fast inference speed is important for models used in dynamic object manipulation.
TABLE III: Analysis on the appearance noise annealing
@@
GEM-N 0.91 (0.87) 0.70 (0.24) 0.86 (0.77) 0.65 (0.16)
GEM-S 0.86 (0.69) 0.84 (0.63) 0.91 (0.74) 0.83 (0.60)
GEM-O 0.93 (0.86) 0.89 (0.83) 0.95 (0.89) 0.87 (0.72)
V-C Component Analysis
@@
This subsection ablates the designs in GEM. When reporting the results, the four evaluation settings described in Section V-A, including simulated in-domain, simulated out-of-domain, real-world tuned sim-to-real, and real-world zero-shot sim-to-real, are denoted as Sim_in, Sim_out, Real_tune, and Real_zero, respectively. Due to the space limit, when reporting the manipulation success rates of these four settings, we report the average success rates of all skills.
V-C1 Analysis on Appearance Noise Annealing
This part analyzes how the appearance noise annealing strategy improves the sim-to-real generalization of GEM. As described in Section IV-B, we apply strong color perturbation to 3D points at the beginning of training and gradually remove the noise as training progresses. To study this strategy, we compare its performance against two baselines, one trained without any color perturbation and another trained with strong color perturbation applied throughout the entire training process. The experimental results are presented in Table III. The original GEM implementation, the GEM without color perturbation, and the GEM with strong color perturbation all the time are denoted as GEM-O, GEM-N, and GEM-S, respectively. The first number in every table cell is the average success rate of four skills, and the second number in a bracket represents the success rate of the Insert skill. We separately list the success rates of the Insert skill because this skill requires highly precise operation and the numerical result change is obvious.
Comparing GEM-O and GEM-N, we can observe that the appearance noise annealing strategy boosts the generalization of the GEM network in unseen environments significantly. For example, in the zero-shot sim-to-real setting of real-robot experiments, the average success rate over different skills is improved by 22%. According to the results of GEM-O and GEM-S, gradually removing the color perturbation achieves better performance than constantly applying strong color perturbation. This is because the visual appearance information is useful in scenarios like distinguishing two closely placed objects. For example, in the Insert task, the network needs to discriminate the inserted object and container, and solely relying on object geometry is difficult to meet this requirement. Therefore, we can find from Table III that the Pick success rates of GEM-S are significantly weaker than GEM-O.
TABLE IV: Comparison between point and image observations.